OpenStudy (anonymous):

A Polyhedron has 10 faces and 16 edges. How many vertices does it have.

OpenStudy (espex):

F + V - E = 2

OpenStudy (anonymous):

You have to use Euler's Method which is F+V-E=2. f=faces, e=edges, and v=vertices so we have to rearrange the formula and solve for vertices since the faces and edges are given so eq becomes V=2+E-F V=2+16-10 V=8